OpenStudy (anonymous):

The monthly charges of two power companies supplying electricity to homes are given by these two equations, where y is the monthly charge in dollars for consumption of x units of electricity. Company A: y = 10 + 2x2 Company B: y = 2.5x At what usage (in number of units) will the monthly charges for both companies be equal? 12.5 units and 50.07 units 5 units and 20.2 units 12.5 units and 20.2 units 5 units and 50.07 units The monthly charges will never be equal.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

hey

jimthompson5910 (jim_thompson5910):

The basic idea is to plug one of the equations into the other. Then solve for x. y = 10 + 2x^2 2.5x = 10 + 2x^2 0 = 10 + 2x^2 - 2.5x 2x^2 - 2.5x + 10 = 0 now use the quadratic formula to solve for x. What do you get when you do so?

OpenStudy (anonymous):

I am getting none of the answers

jimthompson5910 (jim_thompson5910):

what results are you getting?

OpenStudy (anonymous):

this formula?

jimthompson5910 (jim_thompson5910):

correct, you use that formula

OpenStudy (anonymous):

a=2

OpenStudy (anonymous):

b=-2.5

rvc (rvc):

correct :) b=? c=?

OpenStudy (anonymous):

c=10?

jimthompson5910 (jim_thompson5910):

yep, you then plug those values into the formula

rvc (rvc):

yes now try :)

OpenStudy (anonymous):

x=-2.5±sqrt(2.5^2-4(2)(10))/2(2)?

rvc (rvc):

yea

OpenStudy (anonymous):

\[x=-2.5\pm \frac{ \sqrt{2.5^2-4(2)(10)} }{ 2(2) }\]

OpenStudy (anonymous):

I hope that is it

jimthompson5910 (jim_thompson5910):

focus on the stuff inside the square root: 2.5^2-4(2)(10)

jimthompson5910 (jim_thompson5910):

what is that equal to?

OpenStudy (anonymous):

-73.75

rvc (rvc):

solve the equation

jimthompson5910 (jim_thompson5910):

because that value is negative, this means that there are no real solutions for 2x^2 - 2.5x + 10 = 0. The two solutions are complex numbers. So that means the two graphs do NOT intersect at all telling us that the charges will never be equal no matter what x is

OpenStudy (anonymous):

I knew it was something like that

rvc (rvc):

Cool :) complex no here

OpenStudy (anonymous):

:)

rvc (rvc):

All The best!

OpenStudy (anonymous):

Yea I did get a complex number

OpenStudy (anonymous):

didn't think it was what you all was looking for